OpenStudy (anonymous):

The price of an item is initially $900. For the first 4 years the price increases by $35 each year. For the next 4 years the price decreases by $20 per year. From then on the price increases by $35 per year. Find a piecewise-defined function P that specifies the price of the item after t years have passed. Let P(t) = the price of the item in dollars after t years have passed. (Note: Do not use $ signs within your answers.) (b) When 4 < t < or = 8 the value of P(t) is given by what part of the piecewise-defined function? P(t) = ? (c) When t > 8 the value of P(t) is given by what part of the

OpenStudy (anonymous):

sorry just finish the rest of (c) with the same as (b)

OpenStudy (anonymous):

for a) @ 0<=t <= 4 P(t) = 900+35t @ 5<=t<= 8 P(t) = 1040-20t @ t>=9 P(t) = 960 + 35t

OpenStudy (anonymous):

hmmmm that didnt come out as i typed it

OpenStudy (anonymous):

0 < or = t < or = 4 5 < or = t < or = 8 t > or = 9

OpenStudy (anonymous):

that < is the less than sign and > is the greater than

OpenStudy (anonymous):

but all the equations are right. so for b) its given by the second equation of P(t) = 1040 - 20t and c) is the 960 + 35t